OpenStudy (anonymous):

An error exists in the logic that says the following. Explain where and why the error occurs and provide graphical support for your explanation.

OpenStudy (anonymous):

\[\int\limits \tan x dx = \int\limits \frac{ \sin x }{ \cos x } dx\]

OpenStudy (zehanz):

Set u =cos x, then du = -sinx dx, so \(\int \dfrac{\sin x}{\cos x}dx=-\int \dfrac{-\sin xdx}{\cos x}=-\int \dfrac{du}{u}\). Do you recognise what to do now?

OpenStudy (anonymous):

i know i even get the answer \[\ln (\sec) + C\] But I can't find any error

OpenStudy (anonymous):

read the question it says whts the error in converting tan to sin/cos

OpenStudy (zehanz):

Well, if I go on with what I started, I get -ln|u| +C. If u > 0, this becomes \(-\ln u + C = \ln u^{-1} + C = \ln \frac{1}{u}+C=\ln(\sec x) + C\). If u < 0, you have \(-\ln(-u)+C=...=\ln(-\sec x)+C\) So it is necessary to take into cosideration what sign tan x has.

OpenStudy (anonymous):

we need calculus teachers here...not the correct answer

OpenStudy (p0sitr0n):

no brackets should be integral (sinx/cosx)dx

OpenStudy (anonymous):

what

OpenStudy (p0sitr0n):

\[\int\limits tanx dx = \int\limits (\frac{ sinx }{ cosx } )dx\]

OpenStudy (anonymous):

cmon this is not the answer this doesn't even look like an error dude

OpenStudy (p0sitr0n):

lol just guessing maybe its this

OpenStudy (anonymous):

dude i m frikin out lol

OpenStudy (fibonaccichick666):

I am unable to find a flaw in the logic only the math, you should end up with \[ln|cosx|+C\] which you do after integration, I would assume this method approprite

OpenStudy (fibonaccichick666):

correction \[-ln|cosx|+C\]

OpenStudy (anonymous):

ya so whats the error

OpenStudy (fibonaccichick666):

Whenever you do the u substitution integration technique it follows: let u=cosx du=-sinx we then have\[-\int\limits_{}^{}\frac{ 1 }{ u }du\] simply integrate from here the back-substitute yielding \[-ln|u|+C\] then \[-ln|cos x|+C\]

OpenStudy (anonymous):

i know i know...i got this answer too...but i can't understand what is the error...read the question

OpenStudy (fibonaccichick666):

I did and there does not exist an error if it supplies a correct answer we use this trick a lot in higher exponential tangent cases-I think though i may be wrong

OpenStudy (anonymous):

it asks for the error in converting tan to sin/cos

OpenStudy (fibonaccichick666):

yes I understand but as I said I use this alot, I do not see the flaw in the logic provided the cosine function does not equal zero on your bounds

OpenStudy (anonymous):

ok i guess...i said the same thing to my teacher but she said no there is one

OpenStudy (anonymous):

haha

OpenStudy (fibonaccichick666):

ha, really? that is odd. did you google it?

OpenStudy (anonymous):

ya no help from there...google told me to come here haha

OpenStudy (fibonaccichick666):

lol ok well the only issue would be the graph contains asymptotes and DNE at multiple points aka cos x=0 but your domain should be restricted to allow for this. That would be my best guess as to the answer they are fishing for

OpenStudy (anonymous):

what lol...say it again confused

OpenStudy (fibonaccichick666):

h/o a minute

OpenStudy (anonymous):

ok ok

OpenStudy (fibonaccichick666):

so if you look at the graph of the tangent function, it has horizontal asymptotes at every point where cos x=0 (shocking I know). Because cosine has a continuous domain and tangent does not, this must be accounted for whenever you evaluate a definite integral;however, your domain in a tangent function integration, usually accounts for this to start with-unless they are being...not nice people.

OpenStudy (anonymous):

lol...no idea to this man.....

OpenStudy (anonymous):

haha...i am never gonna figure it out XD

OpenStudy (anonymous):

I will copy it and show it to my lecturer and will get back to you

OpenStudy (anonymous):

sweet...dont forget to get the equation from comment box

OpenStudy (anonymous):

yeah, I copied it

OpenStudy (anonymous):

thx tell me ASAP plz

OpenStudy (anonymous):

@amistre64 can u solve this

OpenStudy (amistre64):

i cannot see a reason why there would be an error in that. By definition: \[\tan x=\frac{\sin x}{\cos x}\] is a trig identity. They have the same domains so reiterating the domain as Fibonacci pointed out ... seems moot. There is no error that can determine

OpenStudy (amistre64):

.... that i can determine

OpenStudy (anonymous):

haha...but there is an error idk wht

OpenStudy (zehanz):

@Best_Mathematician: A wide variety of Openstudy members have answered your question. The error you mention cannot be pointed out by any of them. I think it's your turn now. Just saying: there is an error is not enough. Ask your teacher what he/she means with the error. BTW: I am speaking out of experience: sometimes even teachers are wrong :) I know, because I'm a teacher myself...

OpenStudy (inkyvoyd):

are you sure it does not say int tan x dx=int sin x dx/int cos x dx? Because that would be obviously flawed.

OpenStudy (fibonaccichick666):

@inkyvoyd ...that would make so much more sense

OpenStudy (anonymous):

i m heck sure...the question i wrote is the exact same copy...and if it wud be like tht it wud be easy not challenging...and here i m asking challenging questions

OpenStudy (fibonaccichick666):

did you ask your instructor?

OpenStudy (inkyvoyd):

The error is that "an error exists" itself then, I believe. Please ask your instructor :)

OpenStudy (anonymous):

no worry... i will post the answer after my instructor replies...its spring break over here....but still i will reply the answer as soon as i can

OpenStudy (anonymous):

just watch this question

OpenStudy (anonymous):

There is no error in the OP. int(tanx)=int(sinx/cosx)=lnabs(secx)

OpenStudy (anonymous):

maby your getting -ln(abs(cosx)) but using the properties of logarithms this equals ln(abs(cos^(-1)x)) = ln(abs(secx))

OpenStudy (anonymous):

wrong answer

OpenStudy (anonymous):

will reply the correct one asap

OpenStudy (amistre64):

if anything, an error might exist in the assumption that an indefinite integration has a definite answer. \[\int tan(x)~dx=xxx+K\]\[\int \frac{sin(x)}{cos(x)}~dx=xxx+C\] but that distinction would be made pointless by the fact that an indefinite integral represents a family, or set, of functions anyway.